OpenStudy (anonymous):

fan for answer what is the solution the the system 7x+5y=19 -7x-2y=-16 A.(2,1) B.(1,2) C.(7,5) D.(-7,-2) A

OpenStudy (anonymous):

Simplifying 7x + -5y = 19 Solving 7x + -5y = 19 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5y' to each side of the equation. 7x + -5y + 5y = 19 + 5y Combine like terms: -5y + 5y = 0 7x + 0 = 19 + 5y 7x = 19 + 5y Divide each side by '7'. x = 2.714285714 + 0.7142857143y Simplifying x = 2.714285714 + 0.7142857143y

OpenStudy (misty1212):

lol HI!!

OpenStudy (anonymous):

Answer would be A. (2,1)

OpenStudy (misty1212):

\[7x+5y=19 \\-7x-2y=-16 \] add them up in a column and get \[3y=3\]

OpenStudy (misty1212):

that tells you \(y=1\)

OpenStudy (anonymous):

thanksss

OpenStudy (misty1212):

yw \(\color\magenta\heartsuit\)

OpenStudy (anonymous):

Umm...

OpenStudy (anonymous):

could you help me with a few more??

OpenStudy (anonymous):

Me? Yeah..

OpenStudy (anonymous):

Thank you thank you thank you i really suck at math

OpenStudy (anonymous):

It's cool, post your question.. please...or open a new tab and tag me, k?

OpenStudy (anonymous):

which method would be the simplest way to solve the system y=1/2x 2x=3y=28 graphing- my choice substitution elimination distributive

OpenStudy (anonymous):

Substitution. Replace y with x/2 in the second equation and you can quickly solve for x. I don't think they are expecting you to solve the system but just which method is the simplest. But if they want a solution then: Put y = x/2 in second equation: 2x + 3x/2 = 28 multiply throughout by 2 4x + 3x = 56 7x = 56 x = 8 y = x/2 = 8/2 = 4 x = 8, y = 4

OpenStudy (anonymous):

@anti-mathGirl

OpenStudy (anonymous):

Ohh i get it @im.celibate

OpenStudy (anonymous):

Yep, do you need anymore help?

OpenStudy (anonymous):

not now but i'll let you know if i do

OpenStudy (anonymous):

@im.celibate